OpenStudy (anonymous):

65% of 80 is what number? A) 25 B) 52 C) 81.3 D) 123

OpenStudy (anonymous):

\[0.65 * 80 = 52 \]

OpenStudy (anonymous):

Yup Its B. 52.... 80 - 65% = 28,, 28 + 52 = 80...there for its 52